# Current Cross-Correlation Spectroscopy of Majorana Bound States

### v1

@@ Abstract
The clock speed of topological quantum computers based on Majorana zero mode (MZM)-supporting nanoscale devices is determined by the time taken for electrons to traverse the device. We employ the time-dependent Landauer-Büttiker transport theory for current cross-lead correlations in a superconducting nanowire junction hosting MZMs. From the time-dependent quantum noise, we are able to extract traversal times for electrons crossing the system. After demonstrating a linear scaling of traversal times with nanowire length, we present a heuristic formula for the traversal times which accurately captures their behaviour. We then connect our framework to a proposed experimental verification of this discriminant between spurious and genuine MZMs utilizing time-resolved transport measurements.
Introduction.— Modern topological computing protocols are increasingly dominated by the exchange and manipulation of non-local, zero-energy, zero-charge, zero-spin quasiparticles known as Majorana Zero Modes (MZMs) [1, 2, 3]. Recent experimental work has demonstrated the ability to engineer MZMs in superconducting heterostructures, particularly those based on InAs or InSb [4, 5, 6, 7], and more recently Ge [8].
@@
In this work, we use dynamical current correlations and electron traversal times [27, 28] to identify Majorana zero modes (see Fig. 1(b)). Earlier studies have explored noise in MZM-supporting nanowires [29, 30, 31, 32], and high-frequency noise and cross-correlation measurements enable time-resolved detection in mesoscopic conductors [33, 34, 35, 36, 37, 38, 39, 40]. Motivated by the noninvasive Fano-factor approach [32], we show that the nonequilibrium current cross-correlation yields a parameter-tolerant signature present for topological MZMs and absent for trivial states. To our knowledge, these transient Majorana signatures have not been identified previously.
The nonequilibrium Green’s function methodology [41] is well-suited to the study of electron transport in superconducting nanowires [42, 13, 43, 44, 45, 28, 46]. Recent years have seen the development of various approaches to time-dependent quantum transport based on the Green’s function methodology, including the time-dependent Meir-Wingreen [47, 48, 49] and Landauer-Büttiker (TDLB) approaches [50, 51, 52], as well as the time-linear generalized Kadanoff-Baym ansatz for modeling correlated dynamics [53, 54, 55, 56, 57]. Notably, the TDLB theory is a single-shot approach where time steps can be evaluated independently of each other, which enables trivial parallelization, making it computationally more attractive even compared to the time-linear theory.
@@
Since $\mathrm{Re}[C_{LR}(t+\tau,t)]=\frac{1}{2}\langle\{\Delta\hat{I}_{L}(t+\tau),\Delta\hat{I}_{R}(t)\}\rangle$ [61], $P$ coincides with the Blanter-Büttiker cross correlation in reciprocal, stationary settings ($\mathrm{Re}\,C_{LR}=\mathrm{Re}\,C_{RL}$) [62]. Also in non-reciprocal cases, Eq. (2) provides a lead-exchange symmetric measure of cross correlation. Under stationarity, all quantities depend only on $\tau$, and the associated Fourier transformation, $P(\omega)$, is typically referred to as noise power.
Unlike a dc Fano factor, which time-averages over local and nonlocal channels and can be mimicked by trivial (quasi-)Andreev bound states, transient measurements avoid many of these ambiguities documented in realistic nanowires [63]. In particular, time-resolved cross-lead noise isolates crossed-Andreev conversion and reveals a finite traversal time across the wire, a genuinely nonlocal signature expected when spatially separated Majoranas dominate transport [29, 64]. Consistently, finite-frequency or waiting-time diagnostics predict qualitative differences between topological and trivial cases, reinforcing the advantage of the time domain as a discriminator [65, 66, 67].
@@
The model.— The non-interacting TDLB picture can be extended to describe the superconducting nanowire in proximity to a bulk $s$-wave superconductor shown in Fig. 1(a). This is a $1D$ nanowire sandwiched between two normal metal conductors, featuring a strong spin-orbit interaction characteristic of, e.g. InSb or Ge. The spin therefore naturally aligns with the $\pm y$-direction, however, a strong external applied $B$-field along the axis of the nanowire breaks time-reversal symmetry and cants the spins along the $\pm z$-direction, introducing a Zeeman splitting $V_{Z}=g\mu_{B}B/2$, where $\mu_{B}$ is the Bohr magneton and $g$ is the Landé factor.
@@
The method.— We generalize the TDLB approach to the quantum noise [60] so that all Green’s functions appearing in the current-current correlation function are given in the Nambu representation, where each of the $N$-sites on the nanowire shown in Fig. 1(a) corresponds to four Nambu states. For inter-site spacing $a$ the wire length $L=Na$. The one-particle Green’s function in the Nambu representation is defined as
@@
where $\mathcal{T}_{KP}$ denotes an ordering of times $z_{1}$, $z_{2}$ along the complex Konstantinov-Perel’ contour [70], and the brackets denote a quantum statistical average. We set $\hbar=e=k_{\mathrm{B}}=1$ throughout this work. Quantities denoted in bold are $4N\times 4N$-dimensional matrices. $\mathbf{G}$ satisfies the Kadanoff-Baym integro-differential equations of motion on the contour [71], with integral kernel given by the embedding self-energy $\left[\mathbf{\Sigma}_{em}\left(z_{1},z_{2}\right)\right]_{mn}=\underset{k\gamma}{\sum}\tilde{T}_{mk\gamma}\left[\mathbf{g}_{\gamma\gamma}\left(z_{1},z_{2}\right)\right]_{kk}\tilde{T}_{k\gamma n}$, where $\mathbf{g}_{\gamma\gamma}$ denotes the ‘bare’ Green’s function of lead $\gamma$ and the $\tilde{T}_{mk\gamma}$ are defined in the lead-nanowire coupling $\hat{H}_{c}$. All self-energy components can be expressed in terms of $\mathbf{\Gamma}_{\gamma}=2\pi\sum_{k}\tilde{T}_{mk\gamma}\tilde{T}_{k\gamma n}\delta\left(\varepsilon_{\gamma}^{F}-\tilde{\varepsilon}_{k\gamma}\right)$, where $\varepsilon_{\gamma}^{F}$ is the Fermi energy of lead $\gamma$ [72]. The level-width $\mathbf{\Gamma}=\sum_{\gamma}\mathbf{\Gamma}_{\gamma}$ describes the energy-independent lead-nanowire coupling within the wide-band limit approximation [73]. This approximation is valid because we aim to study low-energy excitations near the Fermi level and the normal metal leads can be safely assumed to have broad energy bands compared to the Majorana energy scale.
@@
The trace in Eq. (Current Cross-Correlation Spectroscopy of Majorana Bound States) is taken with respect to the nanowire degrees of freedom. We introduce the lead matrices $\mathbf{\Lambda}^{+}$ and $\mathbf{\Lambda}^{-}$, which correspond to electrons and positively charged holes, respectively, propagating from the leads, as $\mathbf{\Lambda}_{\gamma}^{\pm}\left(t_{1},t_{2}\right)=\pm ie^{-i\psi_{\gamma}\left(t_{1},t_{0}\right)}\int\frac{d\omega}{2\pi}f\left(\pm\omega\right)e^{-i\omega\left(t_{1}-t_{0}\right)}\mathbf{\Gamma}_{\gamma}\mathbf{S}_{\gamma}^{\dagger}\left(t_{2};\omega\right)$ [60]. Physically, the terms in Eq. (Current Cross-Correlation Spectroscopy of Majorana Bound States) are given by lead-lead, lead-nanowire and nanowire-nanowire $e$-$h$ interference effects. Refs. [72] and [60] demonstrate explicitly how the standard Landauer-Büttiker formalisms for the current and current fluctuations, respectively, are recovered from the TDLB formalism in the long-time and static bias limits.
Traversal times.— In the case of the two-terminal system shown in Fig. 1(a), the electronic traversal time $\tau_{\mathrm{tr}}$ can be extracted directly from the lead-symmetrized correlator data, shown in Fig. 1(b), as half of the distance separating the first resonant peaks along the $\tau$-axis. This can be defined more explicitly as the value of the relative delay $\tau$ at which the stationary current cross correlation of Eq. (2) reaches its maximum within a restricted domain $\mathcal{T}$:
@@
We take the argmax over the restricted domain $\mathcal{T}=[\tau_{\mathrm{min}},\tau_{\mathrm{max}}]$, chosen to exclude trivial near-zero features (before the clean ballistic transit) and later-time reflections [61]. For a simplified model nanowire, $\mu=\alpha=V_{Z}=\Delta=0$, $\tau_{\mathrm{tr}}(k)=L/v_{g}(k)$ with $L=Na$ and the group velocity $v_{g}(k)=d\omega_{k}/dk=Ja\sin(ka)$. The fastest mode occurs at $ka=\pi/2$ leading to $\tau_{\mathrm{min}}=N/J$. The upper bound may be set, e.g., to $\tau_{\mathrm{max}}=2\tau_{\mathrm{min}}$ to isolate the first through-wire peak.
@@
Results.— We carry out calculations of the time-dependent cross-correlations for four distinct transport regimes, each specified by a choice of the parameters in our model of a nanowire. The parameters $\mu$, $J$, $V_{Z}$ and $\Delta$ defined above are sufficient to characterize the ordinary and topological superconducting regimes. In addition we study the case of magnetic impurities situated at the wire ends which is characterized by the modified tight-binding parameter $\tilde{V}_{Z}$ described in Ref. [43]. The spurious case of the quasi Majorana mode-supporting nanowire is defined by the usual parameters and an additional smooth confining potential within the nanowire (see Eq. (13) in Ref. [43]). All energies are measured with respect to $J$; times and temperatures are given in terms of $J^{-1}$.
The parameters chosen for our calculations are displayed explicitly in Table 1. In these situations, low-energy states arise from different physical mechanisms with distinct phenomenology. In the trivial superconducting phase, the spectrum is fully gapped and any subgap features correspond to conventional Andreev bound states, which are sensitive to local perturbations [92]. By contrast, MZMs in the topological phase appear as spatially separated, bound, exponentially-decaying end states pinned near zero energy, producing robust zero-bias conductance features and exhibiting nonlocal correlations [93]. Magnetic impurities can induce states occasionally crossing zero energy, but these remain strongly localized and tunable only via local magnetic parameters, lacking topological protection [94]. Also, smooth confinement potentials can produce quasi-Majorana states that mimic many spectroscopic signatures of true MZMs, including apparent zero-bias robustness, yet differ in their strong dependence on confinement geometry and absence of nonlocal parity correlations [95]. Only the low-energy in-gap states are excited by applying a sudden bias voltage $V_{L}(t)=-V_{R}(t)=V_{0}\theta(t)$ with $V_{0}=\Delta/2$.
@@
We then compute the cross-correlation $P\left(t+\tau,t\right)$ for each parameter set, and present its absolute value normalized by the square of the lead-nanowire coupling strength $\Gamma=0.01$ defined between the two end-sites of the wire and the lead, for each model regime in Fig. 2. Details of the numerical implementation can be found in Appendix B of Ref. [28]. Crucially for our method, the implemented formula is a ‘single-shot’ function in the two-time plane, so there is an identical cost associated with taking the observation time $t$ to be very small or very large. In practice, we set the observation time to $t=5000$, i.e. we work directly in the steady-state regime. We also consider the low-temperature regime (to avoid thermal suppression of topological features) by setting $\beta=200$. With the hopping energies $J$ in sub-eV range for prototypical semiconductor nanowires [96], this corresponds to cryogenic temperatures $\lesssim 10$ K.
Immediately we see by comparing the case of the true topological superconductor (Fig. 2(b)) to the spurious Majorana case (Fig. 2(d)) that there is a qualitative distinction between the corresponding cross-correlation functions. Specifically, in Fig. 2(d) we observe strong, unique resonant peaks at $\tau\simeq\pm 50$, corresponding to the traversal of a coherent wavefront across the nanowire, followed by an oscillatory peak structure caused by internal reflections. This is very similar to the behaviour observed in Ref. [27] for a graphene nanoribbon junction. However, for the genuine topological superconductor in Fig. 2(b), the initial electron traversal event is strongly suppressed with respect to modes effective on a significantly longer timescale.
@@
Figure 3 shows the traversal times evaluated, in accordance with Eq. (9), for each model regime as a function of the nanowire length $L$. This data clearly supports the argument of a traversal time as the dependence on the wire length is linear. We can roughly decompose the traversal time as $\tau_{\mathrm{tr}}=\tau_{0}+\tau_{\mathrm{cont}}$, where $\tau_{0}$ corresponds to a coherent wavefront traversing the system. The delay time due to the localization close to the contacts $\tau_{\mathrm{cont}}$ is related to the localized edge states and how strongly they hybridize in the presence of leads [61].
These observations can be translated into an heuristic formula for the traversal time if we evaluate it from the sum of the electron dwell times in a site-wise manner, i.e. $\tau_{\mathrm{tr}}=\sum^{N}_{i=1}\tau_{i}$. A site-wise expression for the $\tau_{i}$ (given explicitly in the Supplemental Material [61]) is then evaluated by choosing an effective intra-site coupling $\lambda_{i}$ which follows the inverse of the on-site potential. The resulting heuristic expression is given by
@@
where $\delta$ controls the exponential drop-off at each end of the wire, $z$ is the localization length evaluated to the nearest integer, the on-site velocity is numerically estimated to be $v_{i}=\frac{\pi}{2}\lambda_{i}$ and $N=L/a$ is the wire length. Note that this expression scales linearly with $N$, as expected. The estimated traversal times $\tau_{\mathrm{tr}}\sim 100$ in units of inverse hopping would correspond to a temporal resolution of picoseconds, for prototypical semiconductor nanowires.
Conclusions.— In this work we analyzed time-resolved cross-lead current correlations in proximitized nanowire junctions within the TDLB framework, using a lead-symmetrized correlator to operationally define and discern an electronic traversal time from transient noise. We evaluated this diagnostic across conventional superconducting, topological Majorana, magnetic-impurity, and quasi-Majorana regimes, finding a robust linear growth of traversal time with wire length and regime-specific temporal fingerprints that distinguish genuine MZMs from spurious near-zero modes. We further distilled a compact heuristic capturing both length scaling and end-localization effects, and verified its quantitative accuracy against simulations. Finally, the predicted picosecond temporal scale expected for prototypical semiconductor nanowires indicates that traversal-time echoes should be observable with current experimental platforms.
To further estimate the time scales relevant for experimental observation, Mourik et al. [4] reported an induced superconducting energy gap of about $250$ µeV in a hybrid superconductor-semiconductor nanowire device. This energy scale corresponds to characteristic in-gap oscillations with a period of roughly $16$ ps, or equivalently a frequency of about $60$ GHz. On the other hand, McIver et al. [97] have demonstrated high-frequency transport measurements using photoconductive switches capable of sub-picosecond temporal resolution. Our framework, employing a noninvasive current cross-correlation probe to identify and distinguish such in-gap states, would therefore call for a stroboscopic implementation of a two-lead setup, where time-resolved currents are detected at both contacts. In this configuration, the nanowire could also be driven out of equilibrium optically rather than by the voltage-bias quench considered in this work.

### v2

@@ Abstract
The clock speed of topological quantum computers based on Majorana zero mode (MZM)-supporting nanoscale devices is limited by the time taken for electrons to traverse the device. We employ the time-dependent Landauer-Büttiker transport theory for current cross-lead correlations in a superconducting nanowire junction hosting MZMs. From the time-dependent quantum noise, we are able to extract traversal times for electrons crossing the system. After demonstrating a linear scaling of traversal times with nanowire length, we present a heuristic formula for the traversal times which accurately captures their behaviour. We then connect our framework to a proposed experimental verification of this discriminant between spurious and genuine MZMs utilizing time-resolved transport measurements.
I Introduction
Modern topological computing protocols are increasingly dominated by the exchange and manipulation of non-local, zero-energy, zero-charge, zero-spin quasiparticles known as Majorana Zero Modes (MZMs) [1, 2, 3]. Recent experimental work has demonstrated the ability to engineer MZMs in superconducting heterostructures, particularly those based on InAs or InSb [4, 5, 6, 7], and more recently Ge [8].
@@ I Introduction
In this work, we use dynamical current correlations and electron traversal times [27, 28] to identify Majorana zero modes (see Fig. 1(b)). Earlier studies have explored noise in MZM-supporting nanowires [29, 30, 31, 32, 33, 34], and high-frequency noise and cross-correlation measurements enable time-resolved detection in mesoscopic conductors [35, 36, 37, 38, 39, 40, 41, 42]. Moreover, shot noise and correlation measurements have been proposed as a basis for discriminating genuine MZMs in recent publications [32, 43]. Motivated by the noninvasive Fano-factor approach [32], we show that the nonequilibrium current cross-correlation yields a parameter-tolerant signature present for topological MZMs and absent for trivial states. To our knowledge, these transient Majorana signatures have not been identified previously.
The nonequilibrium Green’s function methodology [44] is well-suited to the study of electron transport in superconducting nanowires [45, 13, 46, 47, 28, 48]. Recent years have seen the development of various approaches to time-dependent quantum transport based on the Green’s function methodology, including the time-dependent Meir-Wingreen [49, 50, 51] and Landauer-Büttiker (TDLB) approaches [52, 53, 54], as well as the time-linear generalized Kadanoff-Baym ansatz for modeling correlated dynamics [55, 56, 57, 58, 59]. Notably, the TDLB theory is a single-shot approach where time steps can be evaluated independently of each other, which enables trivial parallelization, making it computationally more attractive even compared to the time-linear theory.
@@
It is shown explicitly in Appendix A that
$\mathrm{Re}[C_{LR}(t+\tau,t)]=\frac{1}{2}\langle\{\Delta\hat{I}_{L}(t+\tau),\Delta\hat{I}_{R}(t)\}\rangle$ (3)
and therefore $P$ coincides with the Blanter-Büttiker cross correlation in reciprocal, stationary settings ($\mathrm{Re}\,C_{LR}=\mathrm{Re}\,C_{RL}$) [63]. Also in non-reciprocal cases, Eq. (2) provides a lead-exchange symmetric measure of cross correlation. Under stationarity, all quantities depend only on $\tau$, and the associated Fourier transformation, $P(\omega)$, is typically referred to as noise power.
While a dc Fano factor based protocols can provide valuable diagnostics in appropriately designed setups [32, 43], they are often time-averaged and do not directly resolve propagation delays. Here we therefore focus on the full time-domain cross-lead correlator, whose echo structure separates prompt local processes from delayed nonlocal ones. It also enables an operational extraction of the traversal time across the wire — a genuinely nonlocal signature expected when spatially separated Majoranas dominate transport [29, 64]. Consistently, finite-frequency or waiting-time diagnostics predict qualitative differences between topological and trivial cases, reinforcing the advantage of the time domain as a discriminator [65, 66, 67].
@@
II The model
The non-interacting TDLB picture can be extended to describe a proximitized superconducting nanowire, where superconductivity is induced by a bulk $s$-wave superconductor as shown schematically in Fig. 1(a). This is a $1D$ nanowire sandwiched between two normal metal conductors, featuring a strong spin-orbit interaction characteristic of, e.g. InSb or Ge. The spin therefore naturally aligns with the $\pm y$-direction, however, a strong external applied $B$-field along the axis of the nanowire breaks time-reversal symmetry and cants the spins along the $\pm z$-direction, introducing a Zeeman splitting $V_{Z}=g\mu_{B}B/2$, where $\mu_{B}$ is the Bohr magneton and $g$ is the Landé factor.
@@ II The model
It is worth pointing out that in the present model the pairing field $\Delta$ in $\hat{H}_{w}$ is an effective, static parameter describing proximity-induced superconductivity. The underlying bulk $s$-wave superconductor is therefore not included as an explicit transport terminal. Consequently, the only reservoirs that drive the dynamics and define the current operators are the two normal leads $L$ and $R$, and throughout we focus on their currents and the corresponding cross correlations. We also neglect charging-energy physics and interactions more generally [70, 71, 72]. Within our model, the nonzero cross correlations originate from coherent nonlocal scattering processes that connect the two normal contacts through the proximitized wire.
III The method
We generalize the TDLB approach to the quantum noise [62] so that all Green’s functions appearing in the current-current correlation function are given in the Nambu representation, where each of the $N$-sites on the nanowire shown in Fig. 1(a) corresponds to four Nambu states. For inter-site spacing $a$ the wire length $L=Na$. The one-particle Green’s function in the Nambu representation is defined as
@@ III The method
where $\mathcal{T}_{KP}$ denotes an ordering of times $z_{1}$, $z_{2}$ along the complex Konstantinov-Perel’ contour [73], and the brackets denote a quantum statistical average. In the noninteracting case, this contour is equivalent to the standard Keldysh contour [60]; however the vertical branch on the Konstantinov-Perel’ contour allows for a more intuitive treatment of coupled subsystems prior to the quench time $t_{0}$. We set $\hbar=e=k_{\mathrm{B}}=1$ throughout this work. Quantities denoted in bold are $4N\times 4N$-dimensional matrices. $\mathbf{G}$ satisfies the Kadanoff-Baym integro-differential equations of motion on the contour [74], with integral kernel given by the embedding self-energy $\left[\mathbf{\Sigma}_{em}\left(z_{1},z_{2}\right)\right]_{mn}=\underset{k\gamma}{\sum}\tilde{T}_{mk\gamma}\left[\mathbf{g}_{\gamma\gamma}\left(z_{1},z_{2}\right)\right]_{kk}\tilde{T}_{k\gamma n}$, where $\mathbf{g}_{\gamma\gamma}$ denotes the ‘bare’ Green’s function of lead $\gamma$ and the $\tilde{T}_{mk\gamma}$ are defined in the lead-nanowire coupling $\hat{H}_{c}$. All self-energy components can be expressed in terms of $\mathbf{\Gamma}_{\gamma}=2\pi\sum_{k}\tilde{T}_{mk\gamma}\tilde{T}_{k\gamma n}\delta\left(\varepsilon_{\gamma}^{F}-\tilde{\varepsilon}_{k\gamma}\right)$, where $\varepsilon_{\gamma}^{F}$ is the Fermi energy of lead $\gamma$ [75]. The level-width $\mathbf{\Gamma}=\sum_{\gamma}\mathbf{\Gamma}_{\gamma}$ describes the energy-independent lead-nanowire coupling within the wide-band limit approximation [76]. This approximation is valid because we aim to study low-energy excitations near the Fermi level and the normal metal leads can be safely assumed to have broad energy bands compared to the Majorana energy scale.
@@
The trace in Eq. (III) is taken with respect to the nanowire degrees of freedom. We introduce the lead matrices $\mathbf{\Lambda}^{+}$ and $\mathbf{\Lambda}^{-}$, which correspond to electrons and positively charged holes, respectively, propagating from the leads, as $\mathbf{\Lambda}_{\gamma}^{\pm}\left(t_{1},t_{2}\right)=\pm ie^{-i\psi_{\gamma}\left(t_{1},t_{0}\right)}\int\frac{d\omega}{2\pi}f\left(\pm\omega\right)e^{-i\omega\left(t_{1}-t_{0}\right)}\mathbf{\Gamma}_{\gamma}\mathbf{S}_{\gamma}^{\dagger}\left(t_{2};\omega\right)$ [62]. Physically, the terms in Eq. (III) are given by lead-lead, lead-nanowire and nanowire-nanowire $e$-$h$ interference effects. Refs. [75] and [62] demonstrate explicitly how the standard Landauer-Büttiker formalisms for the current and current fluctuations, respectively, are recovered from the TDLB formalism in the long-time and static bias limits.
IV Traversal times
In the case of the two-terminal system shown in Fig. 1(a), the electronic traversal time $\tau_{\mathrm{tr}}$ can be extracted directly from the lead-symmetrized correlator data, shown in Fig. 1(b), as half of the distance separating the first resonant peaks along the $\tau$-axis. This can be defined more explicitly as the value of the relative delay $\tau$ at which the stationary current cross correlation of Eq. (2) reaches its maximum within a restricted domain $\mathcal{T}$:
@@ IV Traversal times
We take the maximum argument over the restricted domain $\mathcal{T}=[\tau_{\mathrm{min}},\tau_{\mathrm{max}}]$, chosen to exclude trivial near-zero features (before the clean ballistic transit) and later-time reflections discussed in Section V. For a simplified model nanowire, $\mu=\alpha=V_{Z}=\Delta=0$, $\tau_{\mathrm{tr}}(k)=L/v_{g}(k)$ with $L=Na$ and the group velocity $v_{g}(k)=d\omega_{k}/dk=Ja\sin(ka)$. The fastest mode occurs at $ka=\pi/2$ leading to $\tau_{\mathrm{min}}=N/J$. The upper bound may be set, e.g., to $\tau_{\mathrm{max}}=2\tau_{\mathrm{min}}$ to isolate the first through-wire peak.
@@
V Results
We carry out calculations of the time-dependent cross-correlations for four distinct transport regimes, each specified by a choice of the parameters in our model of a nanowire. The parameters $\mu$, $J$, $V_{Z}$ and $\Delta$ defined above are sufficient to characterize the ordinary and topological superconducting regimes. In addition we study the case of magnetic impurities situated at the wire ends which is characterized by the modified tight-binding parameter $\tilde{V}_{Z}$ described in Ref. [46]. The spurious case of the quasi Majorana mode-supporting nanowire is defined by the usual parameters and an additional smooth confining potential within the nanowire (see Eq. (13) in Ref. [46]). All energies are measured with respect to $J$; times and temperatures are given in terms of $J^{-1}$.
@@ V Results
The parameters chosen for our calculations are displayed explicitly in Table 1. In these situations, low-energy states arise from different physical mechanisms with distinct phenomenology. The distinct equilibrium energy spectra for the four regimes considered here are displayed for reference in Fig. 5 of Appendix B. In the trivial superconducting phase, the spectrum is fully gapped and any subgap features correspond to conventional Andreev bound states, which are sensitive to local perturbations [95]. By contrast, MZMs in the topological phase appear as spatially separated, bound, exponentially-decaying end states pinned near zero energy, producing robust zero-bias conductance features and exhibiting nonlocal correlations [96]. Magnetic impurities can induce states occasionally crossing zero energy, but these remain strongly localized and tunable only via local magnetic parameters, lacking topological protection [97]. Also, smooth confinement potentials can produce quasi-Majorana states that mimic many spectroscopic signatures of true MZMs, including apparent zero-bias robustness, yet differ in their strong dependence on confinement geometry and absence of nonlocal parity correlations [98]. Only the low-energy in-gap states are excited by applying a sudden bias voltage $V_{L}(t)=-V_{R}(t)=V_{0}\theta(t)$ with $V_{0}=\Delta/2$.
We then compute the cross-correlation $P\left(t+\tau,t\right)$ for each parameter set, and present its absolute value normalized by the square of the lead-nanowire coupling strength $\Gamma=0.01$ defined between the two end-sites of the wire and the lead, for each model regime in Fig. 2. Details of the numerical implementation can be found in Appendix B of Ref. [28]. Crucially for our method, the implemented formula is a ‘single-shot’ function in the two-time plane, so there is an identical cost associated with taking the observation time $t$ to be very small or very large. In practice, we set the observation time to $t=5000$, i.e. we work directly in the steady-state regime. This is demonstrated to be an observation time at which the cross-correlations have comfortably saturated in Appendix C.
We also consider the low-temperature regime (to avoid thermal suppression of topological features) by setting $\beta=200$. With the hopping energies $J$ in sub-eV range for prototypical semiconductor nanowires [99], this corresponds to cryogenic temperatures $\lesssim 10$ K. The temperature and tunneling-rate dependence of the cross-correlations are examined further in Appendix D, confirming that variations in these parameters do not qualitatively alter the cross-correlation picture.
@@
Immediately we see by comparing the case of the true topological superconductor (Fig. 2(b)) to the spurious Majorana case (Fig. 2(d)) that there is a qualitative distinction between the corresponding cross-correlation functions. Specifically, in Fig. 2(d) we observe strong resonant peaks at $\tau\simeq\pm 50$, which we identify as the coherent wavefront peaks, i.e., the earliest traversal signal across the nanowire. The zero value of the cross-correlation signal at times between these peaks corresponds to the fact that information transfer cannot occur between the metal leads on those timescales. These prominent early-time peaks are followed by an oscillatory peak structure at larger $|\tau|$, consistent with repeated passages through a finite system due to internal reflections. This is very similar to the behaviour observed in Ref. [27] for a graphene nanoribbon junction. In contrast, for the genuine topological superconductor in Fig. 2(b), the early-time wavefront peak is strongly suppressed. The dominant peak structure, with periodic repetition, builds up only on a significantly longer timescale. This general picture of the cross-correlation dynamics is confirmed in Appendix E for various nanowire lengths.
Figure 3: Frequency-resolved noise power spectra obtained from the time-dependent current cross correlations of Fig. 2.
In relation to time-dependent current cross correlations of Eq. (2), it is possible to perform a Fourier transform at the stationary state with respect to the relative time coordinate $\tau$,
$P(\omega)=\int_{-\infty}^{\infty}d\tau\,e^{i\omega\tau}P(t+\tau,t)\quad(t\to\infty),$ (11)
to obtain the noise power spectrum [63]. Figure 3 shows the weighted spectrum $\omega|P(\omega)|$ to suppress the generic low-frequency ($1/f$ type) contribution. We observe that the quasi-Majorana state hosts operational frequencies mostly within the superconducting gap $\Delta$, which correspond to Andreev processes. In contrast, the MZM case shows a pronounced peak around $2\Delta$, which would indicate coupling to continuum states (above the gap) or a pair process. The ordinary superconductor and magnetic impurity cases instead show a broad range of operational frequencies.
Figure 4 shows the traversal times evaluated, in accordance with Eq. (10), for each model regime as a function of the nanowire length $L$. This data clearly supports the interpretation of a traversal time, as the dependence on the wire length is linear. We can roughly decompose the traversal time as $\tau_{\mathrm{tr}}=\tau_{0}+\tau_{\mathrm{cont}}$, where $\tau_{0}$ captures the length-dependent wavefront contribution, while $\tau_{\mathrm{cont}}$ accounts for an additional delay associated with the end regions and contact-induced localization. In this interpretation, the suppression of the early-time wavefront peak in the topological regime (Fig. 2(b)) is reflected in a larger effective contact delay $\tau_{\mathrm{cont}}$, whereas in the quasi-Majorana regime the traversal signal is dominated by the length-dependent wavefront peak (Fig. 2(d)). The delay time $\tau_{\mathrm{cont}}$ is related to the localized edge states and how strongly they hybridize in the presence of leads, as discussed in Appendix F.
These observations can be translated into an heuristic formula for the traversal time if we evaluate it from the sum of the electron dwell times in a site-wise manner, i.e.
$\tau_{\mathrm{tr}}=\sum^{N}_{i=1}\tau_{i}.$ (12)
A site-wise expression for the $\tau_{i}$ (given explicitly in Appendix F) is then evaluated by choosing an effective intra-site coupling $\lambda_{i}$ which follows the inverse of the on-site potential. The resulting heuristic expression is given by
@@
where $\delta$ controls the exponential drop-off at each end of the wire, $z$ is the localization length evaluated to the nearest integer, the on-site velocity is numerically estimated to be $v_{i}=\frac{\pi}{2}\lambda_{i}$ and $N=L/a$ is the wire length. Note that this expression scales linearly with $N$, as expected. The estimated traversal times $\tau_{\mathrm{tr}}\sim 100$ in units of inverse hopping would correspond to a temporal resolution of picoseconds, for prototypical semiconductor nanowires. Equation (13) offers a substantial computational speedup for the evaluation of traversal times in MZM-supporting nanowires compared to the more rigorous TDLB methodology presented in Section III.
VI Conclusions
In this work we analyzed time-resolved cross-lead current correlations in proximitized nanowire junctions within the TDLB framework, using a lead-symmetrized correlator to operationally define and discern an electronic traversal time from transient noise. We evaluated this diagnostic across conventional superconducting, topological Majorana, magnetic-impurity, and quasi-Majorana regimes, finding a robust linear growth of traversal time with wire length and regime-specific temporal fingerprints that distinguish genuine MZMs from spurious near-zero modes. We further distilled a compact heuristic capturing both length scaling and end-localization effects, and verified its quantitative accuracy against simulations. Finally, the predicted picosecond temporal scale expected for prototypical semiconductor nanowires indicates that traversal-time echoes should be observable with current experimental platforms.
To further estimate the time scales relevant for experimental observation, Mourik et al. [4] reported an induced superconducting energy gap of about $250$ µeV in a hybrid superconductor-semiconductor nanowire device. This energy scale corresponds to characteristic in-gap oscillations with a period of roughly $16$ ps, or equivalently a frequency of about $60$ GHz. On the other hand, McIver et al. [100] have demonstrated high-frequency transport measurements using photoconductive switches capable of sub-picosecond temporal resolution. These benchmarks suggest that the time scales highlighted by our calculations set a challenging but potentially accessible target for ultrafast transport experiments. However, we emphasize that realistic measurements will likely be substantially less clean than our idealized simulations, and that implementing time-resolved cross-correlations in a proximitized nanowire is experimentally demanding. We therefore view the time-domain echo structure identified here as a complementary, non-invasive handle on nonlocal transport, rather than a definitive standalone discriminator in all experimental circumstances. In practice, a possible route would be a stroboscopic two-lead implementation where time-resolved currents are detected at both contacts and the device is driven out of equilibrium in a controlled manner. While we focus on a voltage-bias quench in this work, the same framework can accommodate other driving protocols, including optical excitation.